OpenStudy (anonymous):

y=log base 3(x) how would I graph this? is there a way to solve this to graph it?

OpenStudy (anonymous):

the equation you posted is pretty basic and to graph it, you would normally have to give an x and find y then plot that point... and repeat the process with a different x value. a better way is to do an x-y plot of y = 3^x. for example.... when x = 0, y = 1; x = 1, y = 3; x = 2, y = 9; .... you'll get the ordered pairs (0, 1); (1, 3); (2, 9), ... these will plot out y=3^x. to get y=log[base3]x, all you do is reverse the xy of those ordered pairs and plot them.